### v1

# Code: https://github.com/icey-zhang/EfficientMFD

Abstract
[…]
Keywords:
Multimodal images Image Fusion Object Detection
1 Introduction
Precise and dependable object parsing plays a fundamental role in autonomous driving, and remote sensing monitoring [6]. However, merely using visible sensors may fail to accurately recognize objects in complex environments, e.g., inclement weather. Visible-infrared image fusion (VIF) [39, 55, 52, 54] as a typical common multimodal fusion task addresses the above limitation with the complementary information. Numerous multimodal image fusion approaches [15, 16, 19, 11, 8] have been quickly developed to learn the complete representation and complementary information between different modalities. In CDDFuse [51] and DIDFuse [53], a well-designed VIF network is trained first. Then object detection (OD) network is trained using the VIF network results to validate the fusion performance in the separate optimization. Thanks to the powerful capability of deep neural networks on representation learning across multimodalities, deep model-based multimodal fusion has achieved encouraging performance. However, most of them prioritize generating visually pleasing images over addressing downstream high-level visual tasks, posing an obstacle to better object parsing.
Figure 1: Comparison of EfficientMFD with existing VIF-OD task paradigms. Our method redefines the framework composited by image fusion and object detection sub-network in a parallel structure and it can be optimized in one stage joint way which is synchronous. We improve it with a phylogeny tree design for object-region-pixel feature modeling and task alignment to eliminate the task optimization barrier. Dashed arrows indicate cutting off gradient flow.
Recently, some studies have attempted to design a joint learning method by combining the fusion network and high-level tasks e.g. object detection [38] and segmentation [20, 40]. Multimodal fusion detection (MFD) methods have become a hot research issue that VIF and OD enjoy a symbiotic relationship that benefits both parties. VIF generates more informative fused images that significantly improve the performance of OD, while OD provides VIF with rich object semantic information, as its primary objective is to locate and identify objects within the scene. MFD networks are usually designed in a cascaded structure. Joint optimization [18] adopts an OD network as a constraint to a trained VIF network, and thus forces the VIF network to generate fusion images with easily detected objects. Particularly, Zhao et. al [50] proposed a multimodal fusion detection joint learning method, integrating meta-feature embedding from object detection to further promote fusion performance by generating object semantic features through simulating meta-learning. Despite the great success achieved by previous methods as shown in Figure 1, there still remain critical issues. 1) The dominant optimization methods tend to employ a multi-step paradigm in a progressive joint fashion, which significantly undermines efficiency. 2) This optimization overemphasizes the potential of utilizing information from OD to enhance image fusion and makes it difficult to balance the parameters involved, thus being affected by the local optimal solution of individual task. Therefore, seeking unified features for each task simultaneously is still a tough issue. This paper proposed an end-to-end multimodal fusion detection algorithm, named EfficientMFD. EfficientMFD is constructed by fine-grained detail fusion network and coarse-to-fine diffusion-based object detection network, in which the intrinsic features of either one interact via synchronous joint optimization in a one-stage end-to-end manner. To enhance the compatibility of both the fine-grained detail information and semantic content, we introduce the concept of an object-region-pixel phylogenetic tree and coarse-to-fine diffusion process, aiming to mimic the visual perception needed for VIF and OD task requirements. Moreover, we establish a gradient matrix task-alignment (GMTA) to optimize shared complement components to eliminate the undesired effects of inherent optimization. Hence, the gradient can effectively converge to an optimal point with fusion detection weights. Our main contributions can be distilled as follows: (1) We believe that this is the first proposal of the efficient synchronous joint learning named EfficientMFD, which redefined both image fusion and object detection in a one-stage end-to-end manner. We show that this can substantially improve the performance of both tasks. (2) We propose a gradient matrix task-alignment method to measure and quantify the contributions of image fusion and object detection tasks, which help to better balance the stability of the training process and converge to an optimal point with fusion detection weights. (3) Extensive experiments on image fusion and object detection validate the effectiveness of the proposed approach.
2 Related Work
2.1 Multimodal Fusion Object Detection
Due to the powerful nonlinear fitting capabilities of deep neural networks, deep learning has made significant progress in low-level vision tasks, particularly in image fusion tasks [14, 22, 15, 27, 28, 29, 31]. Early efforts [19, 23, 53, 43, 24] tended to achieve excellent fusion results by adjusting network structures or loss functions, overlooking the fact that image fusion should aim to improve the performance of downstream application tasks. Fusion images with good quality metrics may be suitable for human visual perception but may not be conducive to practical application tasks [40]. Some research has acknowledged this issue, Yuan et al[49] utilized various aligned modalities for improved oriented object detection [46, 44, 45, 47] to address the challenge of cross-modal weakly misalignment in aerial visible-infrared images. Liu et al. [18] proposing a joint learning method, pioneered the exploration of the VIF and OD combination methods. Then the optimization of the loss function for segmentation [40] and detection [38] has been validated to be effective in guiding the generation of fused images. They consider the downstream task network as an additional constraint to assist the VIF network in generating fusion results with clearer objects. Zhao et al. [50] leveraged semantic information from OD features to aid VIF and perform meta-feature embedding to generate meta-features from OD features, which are then used to guide the VIF network in learning pixel-level semantic information. Liu et al. [20] proposed multi-interactive feature learning architecture for image fusion and segmentation by enhancing fine-grained mapping of all the vital information between two tasks, so that the modality or semantic features can be fully mutual-interactive. However, OD considers the semantic understanding of objects, while VIF and segmentation primarily focus on the pixel-level relationship between image pairs. The optimization coupling of detection and fusion tasks becomes more challenging to investigate these complementary differences, from which both fusion and detection can benefit. It is worth mentioning that these visible-infrared multimodal fusion detection methods are usually designed in a cascaded structure with tedious training steps. Researchers lack the adoption of end-to-end architectures, which would enable one-step network inference to generate credible fused images and detection results through a set of network parameters.
2.2 Multi-task Learning
Multi-task learning (MTL) [3, 35] involves the simultaneous learning of multiple tasks through parameter sharing. Prior approaches involve manually crafting the architecture, wherein the bottom layers of a model are shared across tasks [2, 13]. Some approaches tailor the architecture based on task affinity [42], while others utilize techniques such as Neural Architecture Search [1, 2, 37] or routing networks [34] to autonomously discern sharing patterns and determine the architecture. An alternative method typically combines task-specific objectives into a weighted sum [12, 5, 25]. In addition, most approaches (e.g. [7, 17, 21, 30, 48]) aim to mitigate the effects of conflicting or dominating gradients. The approach of explicit gradient modulation [21, 30, 48, 36], has demonstrated superior performance in resolving conflicts between task gradients by substituting conflicting gradients with modified, non-conflicting gradients. Inspired by the above multimodal learning methods, we introduce a gradient matrix task-alignment method to align the orthogonal components contained in image fusion and object detection tasks thereby effectively eliminating the inherent optimization barrier that exists between two tasks.
3 The Proposed Method
3.1 Problem Formulation
VIF-OD task concentrates on generating an image that benefits emphasizing objects with superior visual perception capability. The goal of OD is to find the location and identify the class of each object in an image which can naturally provide rich semantic information along with object location information. Therefore, the motivation of OD-aware VIF is to construct a novel infrared and visible image fusion framework that can benefit from the semantic information and object location information contained in object detection. For this purpose, we suppose a pair of visible image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C_{x}}$ and infrared image $\boldsymbol{y}\in\mathbb{R}^{H\times W\times C_{y}}$. The optimization model is formulated as:
$\min_{\boldsymbol{\theta}_{\mathrm{t}}}\mathcal{L}\left(\boldsymbol{t},\mathcal{N}\left(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\theta}_{\mathrm{t}}\right)\right)$ (1)
where $\boldsymbol{t}$ represents the output of the different task network $\mathcal{N}$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{t}}$. $\mathcal{L}(\cdot)$ is a constraint term to optimize the network. Previous approaches solely design image fusion or object detection networks in a cascaded way, which can only achieve outstanding results for one task. To produce visually appealing fused images alongside accurate object detection results, we jointly integrate the two tasks into a unified goal synchronously which can be rewritten as:
$\boldsymbol{\theta}_{\mathrm{u}},\boldsymbol{\theta}_{\mathrm{d}}=\mathrm{arg}\min\omega\mathcal{L}^{\mathrm{u}}\left(\boldsymbol{u},\Phi\left(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\theta}_{\mathrm{u}}\right)\right)+(1-\omega)\mathcal{L}^{\mathrm{d}}\left(\boldsymbol{d},\Psi\left(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\theta}_{\mathrm{d}}\right)\right)+\mathcal{S}\left(\boldsymbol{\theta}^{\star}\right),$ (2)
where $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}_{\mathrm{u}}^{s}=\boldsymbol{\theta}_{\mathrm{d}}^{s}$, defined as the shared parameters for VIF and OD networks. $\boldsymbol{u}$ and $\boldsymbol{d}$ denote the fused image and detection result, which are produced by the VIF network $\Phi(\cdot)$ and OD network $\Psi(\cdot)$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{u}}$ and $\boldsymbol{\theta}_{\mathrm{d}}$. $w$ is a predefined weighting factor to balance the task training. $\mathcal{S}(\cdot)$ is a constrained term to jointly optimize the two tasks. In this paper, we regard the $\mathcal{S}(\cdot)$ as a feature learning constrained manner and achieve this goal by designing a gradient matrix task-alignment training scheme.
Figure 2: An overview of the proposed EfficientMFD framework, which consists of a backbone, nodes, and branches. The backbone is utilized to extract multimodal image features. A fine-grained fusion network (ORPPT) and diffusion-based object detection network (CFDP) are optimized by synchronous joint optimization in an end-to-end manner. Branches include one pixel feature mining module(PFMM) and $L$ region feature refined module (RFRM) to mine different-grained region representations. The gradient domain and conflict between the two tasks are eliminated by the gradient matrix task alignment.
3.2 Architecture
Our proposed EfficientMFD is designed with parallel principle, composited by image fusion and object detection sub-network. Details of the whole architecture are shown in Figure 2. A fusion network (ORPPT) and object detection network (CFDP) can sufficiently realize the granularity-aware detail information and semantic information extraction. Object-Region-Pixel Phylogenetic Tree. The important fact of that humans pay more attention to different regions from coarse to fine for object detection in object scale and image fusion in pixel scale. Inspired by a phylogenetic tree, to simulate humans to study the interactions of hierarchies under different granularity views, we construct an Object-Region-Pixel Phylogenetic Tree (ORPPT) as $\Phi(\cdot)$ to extract different features in multiple region scales. Given an image pair $\boldsymbol{x}\in\mathbb{R}^{W\times H\times C_{x}}$, $\boldsymbol{y}\in\mathbb{R}^{W\times H\times C_{y}}$, we firstly extract image features $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ by shared parallel backbone to save memory and computing resources. Then these features are added in channel dimension to obtain the final $L$ multimodal image features $\boldsymbol{o_{1}},...,\boldsymbol{o_{l}},...,\boldsymbol{o_{L}}\in\mathbb{R}^{W_{1}\times H_{1}\times C_{1}}$. The parameters in $f(\cdot)$ are shared with the OD network. Although $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ can describe the characteristics of visible and infrared modalities, they lack insights into the multi-granularity perspective. Therefore, we utilize the branches including the one pixel feature mining module (PFMM) and $L$ region feature refine module (RFRM) to mine the multiple granularities from coarse to fine. The PFMM $B_{0}$ is set the same with feature fusion block in the MetaFusion [50] with the input pair $\boldsymbol{x},\boldsymbol{y}$. We set $1,2,…,l,…L$ to denote each region branch. For branch $l$, a CNN $\varphi_{l}(\cdot)$ is firstly utilized to extract the granularity-wise feature $\varphi_{l}(\boldsymbol{o_{l}})\in\mathbb{R}^{W_{2}\times H_{2}\times C_{2}}$ at the region level. A set of learnable region prompts $\boldsymbol{R}_{l}=\left\{\boldsymbol{r}_{l,m}\in\mathbb{R}^{C_{2}}\right\}_{m=1}^{M_{l}}$ is introduced to define the $M_{l}$ different regions of granularity-wise feature, where $r_{l,m}$ denotes the $m^{\mathrm{th}}$ region prompt at branch $l$. Then the feature vector is mapped into the region mask $\boldsymbol{A}_{l}=\left\{\boldsymbol{a}_{l,m}\in\mathbb{R}^{W_{2}\times H_{2}}\right\}_{m=1}^{M_{l}}$ by conducting the dot product between the feature vector and region prompt followed by batch normalization and a ReLU activation:
$\boldsymbol{A}_{l}=\mathrm{ReLU}(\mathrm{BN}(\boldsymbol{R}_{l}\cdot\varphi_{l}(\boldsymbol{o_{l}})).$ (3)
Finally, the vector of the object-level feature is weighted by the region mask and further aggregated to form region representation:
$b_{l,m}^{i,j}(\boldsymbol{o_{l}})=\boldsymbol{a}_{l,m}^{i,j}\varphi_{l}^{i,j}(\boldsymbol{o_{l}}),$ (4)
where $b_{l,m}(\boldsymbol{o_{l}})$ denotes the $m^{\mathrm{th}}$ region representation and $(i,j)$ denotes the spatial location. These region-level representations are further concatenated to form the observation $B_{l}(\boldsymbol{o_{l}})=\left[b_{l,1}(\boldsymbol{o_{l}}),b_{l,2}(\boldsymbol{o_{l}}),\ldots,b_{l,M_{l}}(\boldsymbol{o_{l}})\right]$ of branch $l$. These multi-grained attentions concentrate operation on the spatial location information and the extent of the regions which are similar to the task requirements. The $B_{1}$, $B_{2}$, …, $B_{L}$ are up-sampled to keep the consistent spatial size with the pixel-level fusion features $B_{0}$. Then, the region-level fusion features $B_{1},B_{2},...,B_{L}$ are assembled by addition operation followed by a Convolution with $1\times 1$ kernel + ReLU to reduce the channel numbers. Highlighted region features are extracted by a Convolution with $1\times 1$ kernel + Sigmoid and then injected into pixel-level features by multiplication and addition. Finally, five Convolutions with $3\times 3$ kernel + ReLU layers are constructed to reconstruct the fusion result $u$. Coarse-to-Fine Diffusion Process. Diffusion models, inspired by nonequilibrium thermodynamics, are a class of likelihood-based models. DiffusionDet [4] is the first neural network model to utilize the diffusion model for object detection, introducing a novel paradigm that achieves promising results compared to traditional object detection models. Combining diffusion simulation with the diffusion and recovery process of the object box, the Coarse-to-Fine Diffusion Process (CFDP) introduces it as an efficient detection head to assist fusion networks to focus more on object areas. CFDP model defines a Markovian chain of diffusion forward process by gradually adding noise to a set of bounding boxes. The forward noise process is defined as:
$q\left(\boldsymbol{z}_{t}\mid\boldsymbol{z}_{0}\right)=\mathcal{N}\left(\boldsymbol{z}_{t}\mid\sqrt{\bar{\alpha}_{t}}\boldsymbol{z}_{0},\left(1-\bar{\alpha}_{t}\right)\boldsymbol{I}\right),$ (5)
which transforms bounding boxes $z_{0}\in\mathbb{R}^{N\times 4}$ to a latent noisy bounding boxes $\boldsymbol{z}_{t}$ for $t\in\{0,1,\ldots,T\}$ by adding noise to $\boldsymbol{z}_{0}$. $\bar{\alpha}_{t}:=$ $\prod_{s=0}^{t}\alpha_{s}=\prod_{s=0}^{t}\left(1-\beta_{s}\right)$ and $\beta_{s}$ represents the noise variance schedule. During training stage, a neural network $\Psi_{\boldsymbol{\theta}_{\mathrm{d}}}\left(z_{t},t\right)$ is trained to predict $z_{0}$ from $z_{t}$ by minimizing the training objective with $\ell_{2}$ loss:
[…]
At inference stage, bounding boxes $z_{0}$ is reconstructed from noise $z_{T}$ with the model $\Psi_{\boldsymbol{\theta}_{\mathrm{d}}}$ and updating rule in an iterative way, i.e., $\boldsymbol{z}_{T}\rightarrow\boldsymbol{z}_{T-\Delta}\rightarrow\ldots\rightarrow\boldsymbol{z}_{0}$. In this work, we aim to solve the object detection task via the diffusion model. A neural network $\Psi_{\boldsymbol{\theta}_{\mathrm{d}}}\left(\boldsymbol{z}_{t},t,\boldsymbol{x,y}\right)$ is trained to predict $z_{0}$ from noisy boxes $\boldsymbol{z}_{t}$, conditioned on the corresponding image pair $\boldsymbol{x,y}$.
3.3 Loss Function
The total loss function is combined with an image fusion loss function $\mathcal{L}_{\mathrm{f}}$ and object detection loss $\mathcal{L}_{\mathrm{d}}$. $\mathcal{L}_{\mathrm{f}}$ consists of three types of losses, i.e., structure loss $\mathcal{L}_{\mathrm{SSIM}}$, pixel loss $\mathcal{L}_{\text{pixel }}$ and gradient loss $\mathcal{L}_{\text{grad }}$. For one fused image, it should preserve overall structures and maintain a similar intensity distribution from source images. To this end, the structural similarity index (SSIM) is introduced in function:
$\mathcal{L}_{\text{SSIM }}=\left(1-\operatorname{SSIM}({\boldsymbol{u},\boldsymbol{x}})\right)/2+\left(1-\operatorname{SSIM}({\boldsymbol{u},\boldsymbol{y}})\right)/2,$ (7)
[…]
$\begin{gathered}\mathcal{L}_{\text{pixel}}^{\text{o}}=\left\|I_{m}\circ\left(\boldsymbol{u}-max(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\\ \mathcal{L}_{\text{pixel}}^{\text{b}}=\left\|(1-I_{m})\circ\left(\boldsymbol{u}-mean(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\end{gathered}$ (8)
where $\|\cdot\|_{1}$ stands for the $l_{1}$-norm. $\boldsymbol{\omega_{\mathbf{1}}}$ and $\boldsymbol{\omega_{\mathbf{2}}}$ are saliency weight maps calculated by $\boldsymbol{\omega}_{\mathbf{1}}=\boldsymbol{S}_{\boldsymbol{x}}(k)/\left[\boldsymbol{S}_{\boldsymbol{x}}(k)-\boldsymbol{S}_{\boldsymbol{y}}(k)\right],\boldsymbol{\omega}_{\mathbf{2}}=1-\boldsymbol{\omega}_{\mathbf{1}}$. Supposing that the saliency value of $\boldsymbol{x}$ at the $k^{\mathrm{th}}$ pixel can be obtained by $\boldsymbol{S}_{\boldsymbol{x}}(k)=\sum_{i=0}^{255}\boldsymbol{H}_{\boldsymbol{x}}(i)|\boldsymbol{x}(k)-i|$, where $\boldsymbol{x}(k)$ is the value of the $k$th pixel and $\boldsymbol{H}_{\boldsymbol{x}}(i)$ is the histogram of pixel value $i$. The operator $\circ$ denotes the elementwise multiplication, max($\cdot$) denotes the element-wise maximization, and mean($\cdot$) denotes the element-wise average operation. Therefore, the object-aware pixel loss $\mathcal{L}_{\text{pixel }}$ is defined as:
[…]
Besides, gradient information of images always characterizes texture details, thus, we use $\mathcal{L}_{\text{grad }}$ to constrain these textual factors to a multi-scale manner:
$\mathcal{L}_{\text{grad}}=\sum_{k=3,5,7}\left\|\nabla^{k}\boldsymbol{u}-\max\left(\nabla^{k}\boldsymbol{x},\nabla^{k}\boldsymbol{y}\right)\right\|_{2}^{2},$ (10)
where $\nabla$ denotes gradient operators that calculate by $\nabla$=∇ = $\boldsymbol{u}-\mathcal{G}(\boldsymbol{u})$ with combination of different Gauss $(\mathcal{G})$ kernel size $k$. Totally, we obtain $\mathcal{L}_{\mathrm{u}}=\eta_{1}\mathcal{L}_{\mathrm{SSIM}}+\eta_{2}\mathcal{L}_{\text{pixel }}+\eta_{3}\mathcal{L}_{\text{grad }}$.
3.4 Gradient Matrix Task-Alignment
The VIF and OD tasks have distinct optimization objectives. VIF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. An inherent optimization barrier exists between these two tasks. We observe that the prevailing challenges in multi-task learning are arguably task dominance and conflicting gradients. We introduce a gradient matrix Task-Alignment (GMTA) by presenting the condition number to mitigate the undesired effects of optimization barrier in task-shared parameters $\boldsymbol{\theta}^{\star}$ which are supposed to be balanced between the VIF and OD tasks. The individual task gradients of VIF and OD task are calculated by $\boldsymbol{g}_{u}=\nabla_{{\theta}^{\star}}\mathcal{L}_{u}$ and $\boldsymbol{g}_{d}=\nabla_{{\theta}^{\star}}\mathcal{L}_{d}$ in the training optimization process. The gradient matrix can be defined as $\boldsymbol{G}=\left\{\boldsymbol{g}_{u},\boldsymbol{g}_{d}\right\}$. In multi-task optimization, a cumulative gradient $\boldsymbol{g}=\boldsymbol{G}\boldsymbol{w}$ is a linear combination of task gradients and the stability of a linear system is measured by the condition number of its matrix in numerical analysis. Hence the stability of the gradient matrix is equal to the ratio of the maximum and minimum singular values of the corresponding matrix: $\kappa(\boldsymbol{G})=\frac{\sigma_{\max}}{\sigma_{\min}}$. Learning from the Aligned-MTL [37], The condition number is optimal ($\kappa(\boldsymbol{G})=1$) if and only if the gradients are orthogonal and equal in magnitude. Thereby, we consider the feature learning constraint $\mathcal{S}(\boldsymbol{\theta}^{\star})$ can be defined as the following optimization to eliminate instability in the training process:
$\min_{\hat{\boldsymbol{G}}}\|\boldsymbol{G}-\hat{\boldsymbol{G}}\|_{F}^{2}\quad\text{ s.t. }\hat{\boldsymbol{G}}^{\top}\hat{\boldsymbol{G}}=\boldsymbol{I}.$ (11)
We can use a minimum non-zero singular value of an initial gradient matrix $\sigma(\hat{\boldsymbol{G}})=\sigma_{\min}(\boldsymbol{G})>0$ since a condition number defines the system up to scale. The problem Eq. 11 can be treated as a Procrustes problem and can be solved by performing a singular value decomposition (SVD) to $\boldsymbol{G}$ ($\boldsymbol{G}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T}$) and rescaling singular values corresponding to principal components so that they are equal to the smallest singular value:
$\hat{\boldsymbol{G}}=\sigma\boldsymbol{U}\boldsymbol{V}^{\top}=\sigma\boldsymbol{G}\boldsymbol{V}\boldsymbol{\Sigma}^{-1}\boldsymbol{V}^{T}.$ (12)
The final linear system of gradients defined by $\hat{\boldsymbol{G}}$ satisfies the optimal condition in terms of a condition number. Gradient matrix task-alignment handles issues related to the stability of gradients in the two tasks addressing optimization conflicts.
Figure 3: Visualized results of image fusion on M3FD, TNO and RoadScene.
4 Experiments and Analysis
4.1 Implementation Details
We conduct all the experiments on a server with one GeForce RTX 3090 GPU. The pretrained DiffusionDet is used for the initialization of the object detection network. In the training phase, EfficientMFD is optimized by AdamW with a batch size of 1. We set the learning rate to $2.5\times 10^{-5}$ and the weight decay as $41e-$ 4. The default training iteration is only 15,000. To fully demonstrate the effectiveness of our method, we compare seven other state-of-the-art models, including VIF method (CDDFuse [51], DIDFuse [53], U2Fusion [43] PIAFusion [41], and SwinFusion [26]), and VIF-OD joint method (MetaFusion [50], and Tardal [18]). For more analyses about the other VIF-OD datasets and visualization results, see the Appendix.
4.2 Dataset
We conduct experiments on three widely-used visible-infrared image datasets: TNO, RoadScene, and M3FD [18]. TNO and RoadScene are just used to evaluate VIF performance. Besides, M3FD is adopted to evaluate both VIF and OD performance. RoadScene with 37 image pairs and TNO with 42 image pairs are only used for the VIF task in the testing stage. The VIF network is trained by the M3FD dataset. The M3FD dataset [18] which is divided into a training set (2,940 image pairs) and a testing set (1,260 image pairs) and the 300 image pairs in the testing set is used to validate the VIF task. The original image size in the dataset is $1,024\times 768$.
4.3 Metric
[…]
Table 1: Quantitative results of different fusion methods on TNO, RoadScene and M3FD datasets. The model training (Tr.) and test (Te.) time is counted on a NVIDIA GeForce RTX 3090. The best result is highlighted.
Task Method M3FD TNO RoadScene Tr. Time Te. Time
EN MI VIF EN MI VIF EN MI VIF
VIF DIDFuse[53] 6.13 14.65 1.51 6.30 15.30 1.47 6.67 16.65 1.55 3h9m38s 0.096s
U2Fusion[43] 5.66 14.22 1.50 5.78 14.89 1.49 6.25 16.30 1.57 4h8m36s 2.091s
PIAFusion[41] 5.75 13.92 1.59 5.05 13.61 1.36 6.37 16.22 1.58 5h35m20s 0.003s
SwinFusion[26] 5.80 13.83 1.58 6.09 14.28 1.55 6.30 15.93 1.60 3h38m5s 0.044s
CDDFuse[51] 5.77 13.82 1.58 6.21 15.03 1.49 6.54 16.54 1.57 5h59m59s 0.096s
VIF-OD Tardal[18] 5.72 14.68 1.47 5.87 14.99 1.43 6.72 16.98 1.54 5h36m28s 0.093s
Metafusion[50] 6.20 15.19 1.54 6.29 16.03 1.44 6.35 16.76 1.57 6h47m38s 0.002s
EfficientMFD 6.36 15.47 1.65 6.40 16.28 1.6 6.79 17.11 1.69 2h50m32s 0.014s
4.4 Results of Infrared-visible Image Fusion
Qualitative results of different fusion methods are depicted in Figure 3. All the fusion methods can fuse the main features of the infrared and visible images to some extent and we can clearly observe two remarkable advantages of our method. First, the significant characteristics of infrared images can be effectively highlighted by our method. As shown in Figure 3, our M3FD fusion image captures the person riding a motorcycle, and the TNO fusion image clearly represents the person. In comparison with other methods, our method demonstrates high contrast and recognition of the objects, providing not only a visually pleasing experience but also assistance for object detection tasks. Second, our method preserves rich details from the visible images, including color and texture. In Figure 3, our advantages are evident in the fusion images across the three datasets, such as the clear outline of the white car’s rear, information on the cabinet, and words and arrows on the signboard. While retaining a substantial amount of detail, our method maintains a high resolution without introducing blurriness. In contrast, other methods fail to simultaneously achieve these two advantages. Analyzed together with the results in Table 2, we can get our method accomplishes visual friendliness while effectively aiding downstream tasks. Sequentially, we provide quantitative results of different fusion methods in Table 1. Our EfficientMFD generally achieves the largest metric values. In specific, the largest average value on MI proves that our method transfers more considerable information from both two source images. Values of EN reveal that our results contain edged details and the highest contrast between objects and the background. The large VIF shows our fusion results have high-quality visual effects and small distortion compared with the source images. Moreover, our method achieves the fastest training time and only needs about 3 hours to finish the joint learning at one stage which means that faster iterative updates can be done on new datasets. The test time to generate a fused image ranked third.
Table 2: Quantitative results of object detection on M3FD datasets among all the image fusion method + detector (i.e. YOLOv5s). The best result is highlighted.
[…]
V/I Infrared 49.3 67.1 72.9 35.8 43.6 61.6 85.3 55.1
Visible 38.1 69.4 75.5 44.4 44.8 63.2 86.3 55.9
VIF DIDFusion[53] 45.8 68.8 73.6 42.2 43.7 61.5 86.2 56.2
U2Fusion[43] 47.7 70.1 73.2 43.2 44.6 63.9 87.1 57.1
PIAFusion[41] 46.5 69.6 75.1 45.4 44.8 61.7 87.3 57.2
SwinFusion[26] 44.5 68.5 73.3 42.2 44.4 63.5 85.8 56.1
CDDFuse[51] 46.1 69.7 74.2 42.2 44.2 62.7 87.0 56.5
VIF-OD Tardal[18] 49.8 65.4 69.5 46.6 43.7 61.1 86.0 56.0
MetaFusion[50] 48.4 66.7 70.5 49.1 46.4 59.9 86.7 56.8
Ours (YOLOv5s) 51.0 67.9 69.4 50.2 48.7 61.6 87.9 58.1
EfficientMFD 60.1 69.5 81.4 52.2 47.6 72.2 91.8 63.8
4.5 Results of Infrared-visible Object Detection
Figure 4: Visualize results of object detection on M3FD.
To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. We followed the same parameter settings, and the results are shown in Figure 4. The detection results are poor when using only single-modal image inputs, with instances of missed detections, such as the motorcycle and rider next to the car and people on the far right in the image. Almost all fusion methods reduce missed detections and improve confidence by fusing information from both modalities. Through the design of an end-to-end fusion detection synchronous optimization strategy, we obtained fusion images that are visually and detection-friendly, especially for occluded and overlapping objects, as seen in the blue ellipse with the motorcycle and the overlapping people on the far right in the image.
To further assess the quality of visible and infrared fusion images, we conduct a fair comparison between our method and the state-of-the-art (SOTA) methods on YOLOv5s. As shown in Table 2, the SOTA methods demonstrate an improvement in performance compared to single-modal detection, indicating that well-fused images can effectively assist downstream tasks. In contrast, our fusion method achieves better performance, providing effective prompts for object detection tasks at the image level, especially for the objects that need to be detected. With the integration of our method, the detection performance is further enhanced, reaching the highest accuracy for each class. This better demonstrates the advantages of our training paradigm.
Table 3: Effect study of gradient matrix task-alignment by comparing with different optimization methods on M3FD. The best result is highlighted.
[…]
VIF 6.09 14.90 1.48 / /
OD / / / 90.28 62.75
VIF-OD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
VIF-OD (w GMTA) 6.36 15.47 1.65 91.80 63.83
4.6 Ablation Studies
Effect of Gradient Matrix Task Alignment. In Sec. 3.4, we introduce Gradient Matrix Task Alignment, which helps balance the advantages and conflicting gradients of VIF and OD tasks, making the training process more stable and eliminating optimization barriers between the two tasks. To verify its effectiveness, we compared separate optimizations for VIF and OD, as well as joint optimization. Specifically, VIF represents using only $\mathcal{L}_{u}$ to optimize the fusion network, OD represents using only $\mathcal{L}_{d}$ to optimize the object detection network, and VIF-OD represents simultaneous optimization of the fusion and detection networks using both $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ loss functions. The results, as shown in Table 3, indicate that methods incorporating GMTA optimization constraints with shared weights achieve the best results for both VIF and OD. This is because the VIF and OD tasks have distinct optimization objectives. VIF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. Therefore, simply optimizing the entire network with shared loss functions may be influenced by local optimal solutions of individual task, preventing the joint task from achieving the optimal solution. The accuracy of VIF-OD (w/o GMTA) in the table shows a slight decrease compared to separately training the detection network. In contrast, GMTA orthogonalizes the gradients of shared parameters corresponding to the two tasks, allowing the joint network to converge to an optimal point with fusion detection weights.
Table 4: Average results of ablation study of the number of branches in the Object-Region-Pixel Phylogenetic Tree on M3FD dataset. The best result is highlighted.
[…]
0 6.10 15.19 1.54 91.51 63.20
0,1 6.10 15.29 1.54 91.65 62.96
0,1,2 6.19 15.28 1.58 91.73 63.46
0,1,2,3 6.36 15.47 1.65 91.80 63.83
0,1,2,3,4 5.99 14.31 1.40 91.73 63.55
Study of Branches in the Object-Region-Pixel Phylogenetic Tree. We have implemented the Object-Region-Pixel Phylogenetic Tree to explore the hierarchical interactions under different granularity views and extract various features across multiple region scales, as introduced in Sec. 3.2. Here, we investigate the combination of the pixel feature mining module (0) and the region feature refinement module (1,2,3,4). The results are shown in Table 4. It can be observed that with the increase in the number of branches, the fusion network achieves higher image fusion quality and object detection performance. Region features provide the fusion network with multi-level semantic features of objects. However, when higher-level semantic object information is added, the performance of the fusion network declines. This is because the detailed information contained in the deeper layers of the backbone structure shared with the detection network decreases abruptly, which may affect the fusion network’s absorption of these features, thereby influencing pixel-level fusion tasks. This illustrates the importance of balancing the semantic information provided between the detection network and the fusion network with pixel-level information.
Figure 5: Visualization of task dominance and conflicting gradients in the joint learning of OD and VIF tasks.
4.7 Analysis of Gradient Matrix
As described in Section 3.4, the VIF and OD tasks pursue different optimization goals. VIF primarily focuses on capturing pixel-level relationships between image pairs, while OD integrates object semantics within the broader context of diverse scenes. To visualize the task dominance and conflicting gradients, we plot the gradient matrix in the training stage illustrated by Figure 5. We perform a GMTA operation every 1000 iteration loss updates. Blue represents the gradients of shared parameters computed by the object detection loss function, while yellow represents the gradients of shared parameters computed by the modal fusion loss function. During the training process, it can be observed that the gradient values of the object detection task are larger and dominant, while the gradient values of the fusion task are smaller. This may affect the learning process of the fusion task during training. Conversely, the utilization of GMTA effectively mitigates this gradient dominance and conflict, facilitating a balance of shared parameters between VIF and OD.
5 Conclusion
Within this paper, synchronous-parallel task-alignment optimization is proposed to formulate fusion and detection in a harmonious one-stage training process. We introduce a phylogenetic tree structure to simulate these two tasks in different visual perceptions needed for device task requirements. In addition, we align the orthogonal components of the fusion detection linear system of the gradients. By unrolling the model to a well-designed fusion network and diffusion-based detection network, we can generate a visual-friendly result for fusion and object detection in an efficient way without tedious training steps and inherent optimization barriers.
6 Appendix
6.1 More Fusion Visualization Results
More comparisons of infrared-visible image fusion visualization results are depicted in Figure 6, 7 and 8. These fusion results demonstrate the advantages of synchronously optimizing fusion and detection tasks. With minimal training costs, we obtain fusion images that are visually and detection-friendly. Specifically, the fusion images retain significant object information extracted from infrared images, while also preserving detailed information such as texture, color, and background from visible images. Our method effectively combines the strengths of both modalities to enhance the overall performance of the detection.
Figure 6: Visualize results of image fusion on M3FD.
Figure 7: Visualize results of image fusion on RoadScene.
Figure 8: Visualize results of image fusion on TNO.
6.2 More Detection Visualization Results
The visualization of infrared-visible object detection are demonstrated in Figure 9. Our fusion images are noticeably superior due to the effective combination of object information from infrared images and texture details from visible images. This integration enables the object detection network to distinguish between the background and the object clearly. Additionally, for small and occluded objects, the clear edge details assist the network in achieving improved detection results.
Figure 9: Visualize results of object detection on M3FD.